OpenStudy (anonymous):

calculate y' if x(y)^4 +x^2=x+3y please help and show all steps

OpenStudy (anonymous):

\[xy^4+x^2=x+3y\] \[y^4+3xy^3y'+2x=1+3y'\] solve for \[y'\]